OpenStudy (anonymous):

integrate xe^(2x) dx

OpenStudy (anonymous):

\[\int\limits_{}^{}x e^{2x}\]dx

OpenStudy (abb0t):

Are you familiar with u-substitution? You are going to apply that to this problem.

OpenStudy (abb0t):

u = 2x du = xdx

OpenStudy (anonymous):

yes I am, I figured i would need it but was still a little confused

OpenStudy (anonymous):

but the x doesn't cancel out

OpenStudy (anonymous):

\[\int\limits_{}^{}xe^u du/2\]

OpenStudy (abb0t):

notice how you have du = xdx which you DO have. \[\int\limits e^u(xdx) <-- du\]

OpenStudy (anonymous):

wouldn't du = 2dx though?

OpenStudy (abb0t):

Oh, wait, i think you might want to use integration by parts. sorry.

OpenStudy (anonymous):

by using integration by parts let u = x

OpenStudy (anonymous):

I don't think I've learned that yet, hmm. I was trying to find the volume of something using the cylindrical shell method and my integral ended up being \[2\pi \int\limits_{}^{}2x+xe ^{2x} dx\]

OpenStudy (anonymous):

i just realized I should be using disk method since it's rotating about horizontal line, sorry! I'll give you best answer anyway since you were helping

OpenStudy (abb0t):

\[\int\limits fg'dx= fg-\int\limits f'gdx\] u = f(x) v = g(x) du = f'(x)dx dv = g'(x)dx \[\int\limits udv = uv- \int\limits vdu\]